Ammonia is produced by the reaction of nitrogen and hydrogen.
N2(g)+3H2(g)→2NH3(g)Ammonia
How many moles of H2 are needed to react with 2.3 mol of N2?
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2015-10-20T03:58:58-0400
6,9 moles of H2 are needed to react with 2.3 mol of N2
Explanation:
N (H2) = n (N2) * 3
N (H2) = 2,3 * 3 = 6,9 mol